OpenStudy (anonymous):

3. Solve for x: (x - 9)2 = 1

OpenStudy (anonymous):

x-9 = 1 ==>x=10 or x-9 =-1 ==>x=8

OpenStudy (anonymous):

\[x-9=\pm1\]\[x = 9 \pm 1\]

OpenStudy (anonymous):

(x-9)^2=1 x-9=sqrt1 x-9=1, x-9=-1 x=10, x=8

OpenStudy (anonymous):

its completing the square

OpenStudy (anonymous):

i got x=9+_ squrt 163 is that right

jimthompson5910 (jim_thompson5910):

No, the answers are x = 10 or x = 8

OpenStudy (anonymous):

how can u show me plz

jimthompson5910 (jim_thompson5910):

Start with (x-9)^2 = 1 and take the square root of both sides to get x - 9 = +-sqrt(1)

jimthompson5910 (jim_thompson5910):

Break up the right side to get two equations x-9 = 1 or x - 9 = -1 Now solve for x in each equation to get x = 10 or x = 8